# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

### v1

@@ Abstract.
Graphs are essential data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems. Graph-level tasks, which predict properties or classes for the entire graph, are critical for applications, such as molecular property prediction and subgraph counting. Graph Neural Networks (GNNs) have shown promise in these tasks, but their evaluations are often limited to narrow datasets, tasks, and inconsistent experimental setups, restricting their generalizability. To address these limitations, we propose a unified evaluation framework for graph-level GNNs. This framework provides a standardized setting to evaluate GNNs across diverse datasets, various graph tasks (e.g., graph classification and regression), and challenging scenarios, including noisy, imbalanced, and few-shot graphs. Additionally, we propose a novel GNN model with enhanced expressivity and generalization capabilities. Specifically, we enhance the expressivity of GNNs through a $k$-path rooted subgraph approach, enabling the model to effectively count subgraphs (e.g., paths and cycles). Moreover, we introduce a unified graph contrastive learning algorithm for graphs across diverse domains, which adaptively removes unimportant edges to augment graphs, thereby significantly improving generalization performance. Extensive experiments demonstrate that our model achieves superior performance against fourteen effective baselines across twenty-seven graph datasets, establishing it as a robust and generalizable model for graph-level tasks.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
Graphs represent objects as nodes and their relationships as edges, which are key data structures across various domains to model complex interactions (Bonifati et al., 2024; Fan, 2022; Fang et al., 2022), such as social networks, molecular structures, biological systems, etc. Graph-level tasks involve predicting properties or labels of entire graphs, rather than individual nodes or edges. These tasks are crucial in various domains, such as subgraph counting in database management systems (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020b), and protein classification in biology (Morris et al., 2020; Li et al., 2024a). Recently, graph neural networks (Veličković et al., 2017; Hamilton et al., 2017; Liao et al., 2022; Cui et al., 2021; Li et al., 2024c) have emerged as powerful tools for modeling and analyzing graph-structured data. They learn node representations by iteratively aggregating information from their neighbors and obtain graph representations based on node representations for graph-level tasks.
Depending on the technique, current GNNs designed for graph-level tasks can be categorized into five types: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based GNNs. Firstly, node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) learn representations for individual nodes, which are then used to form a graph-level representation by a Readout function, such as averaging the representations of all nodes. Secondly, HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) employ hierarchical pooling techniques to progressively reduce the graph size while preserving its essential structural information. HP-based GNNs aim to capture multi-level graph structures to learn a graph representation.
Thirdly, subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) separate a graph into several subgraphs. They learn the representations for each subgraph and then obtain the final graph representation based on these subgraph representations. Fourthly, GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d; Li et al., 2024b) assume that graphs are noisy and incomplete. These methods aim to reconstruct the graph structure and node features to enhance graph quality, enabling the learning of more accurate and reliable graph representations. Finally, SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) aim to pretrain GNNs on unlabeled graphs to learn generalizable node representations, which can then be fine-tuned with labeled data. Specifically, they pretrain GNNs by either predicting existing graph properties (e.g., node degrees and attributes) (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Wang et al., 2022b) or maximizing the similarity between graph representations obtained from different augmentations (e.g., edge dropping) of the same graph (Lee et al., 2022; Hassani and Khasahmadi, 2020; Li et al., 2024a; Sun et al., 2024).
@@
Although various GNNs being designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2019; Wang et al., 2024a). This limited scope fails to accurately reflect the advancements and capabilities of existing GNNs in handling graph-level tasks. To address the need for fair and comprehensive empirical evaluations, we first identify three key issues in existing evaluation frameworks, thereby enhancing our evaluation capabilities.
@@
  • Issue 1. Insufficient Coverage of Data Domains. The evaluations of existing approaches often concentrate on a limited range of datasets, typically in specific domains such as chemistry or biology. This narrow focus can cause models to overfit to the unique characteristics of these datasets, limiting the generalizability and applicability of GNNs to other domain datasets, such as social networks or geometric graphs.
  • Issue 2. Insufficient Graph-level Tasks and Scenarios. Current evaluation frameworks only focus on one type of graph-level task, such as graph classification for molecular graphs. The insufficient coverage of graph tasks neglects the diversity of potential graph-level applications, such as cycle counting and path counting in graphs. Additionally, current methods only assess GNNs under the assumption of enough labeled graphs. However, there are many other realistic scenarios. For example, the graphs labels are few and imbalanced, and the graph datasets are noisy.
  • Issue 3. Lacking One Unified Evaluation Pipeline. The absence of a standardized evaluation pipeline leads to inconsistent reporting and comparison of results across studies. Different works may use varying metrics, datasets, and experimental setups, complicating the assessment of progress and the identification of truly effective models. A unified evaluation framework would facilitate more transparent and reliable comparisons
@@
To address the above three issues, we propose a comprehensive framework to ensure a more comprehensive and fair assessment of GNNs for graph-level tasks. Firstly, we categorize existing GNNs for graph-level tasks into five types and analyze each type in depth. Secondly, we propose a unified and fair evaluation pipeline with standardized data splitting and experimental settings. We use graph datasets from diverse domains. including biology, chemistry, social networks, and geometric graphs. Thirdly, we evaluate existing graph-level GNNs on both graph classification and graph regression (e.g., cycle counting). We also consider diverse scenarios beyond those with ample labeled graphs, including noisy graphs, imbalanced graphs, and the few-shot setting.
Additionally, alongside the comprehensive evaluation framework, we propose an improved GNN model for graph-level tasks, which achieves both high expressivity and strong generalization capabilities. Firstly, to enhance the expressivity of GNNs, we introduce a general $k$-path rooted subgraph GNN, which samples subgraphs for each $k$-length path starting from every node. Our model learns representations for each subgraph and aggregates these into a comprehensive graph representation. We theoretically prove that k-path rooted subgraph GNNs can, at a minimum, count the number of $k+2$-Path and $k+3$-Cycle. Secondly, to improve generalization across diverse domains and scenarios, we propose an adaptive graph contrastive learning approach to augment graphs. By analyzing edge importance across different domains, we propose a unified metric based on node degree to remove unimportant edges, ensuring the augmented graphs retain their original structural patterns. Our model is then optimized by maximizing the similarity between augmented graphs and predicting labels consistently with the original graph. This approach makes our model resistant to noise, thereby improving its generalization ability.
@@
We summarize the contributions of this paper as follows.
  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, i.e., node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based. We provide an in-depth analysis of their distinct techniques for learning graph representations, offering a clear understanding of their strengths and limitations.
  • We propose a unified evaluation framework to address limited domain diversity, insufficient task coverage, and the lack of standardization. It includes diverse datasets, supports multiple tasks, and ensures fair comparisons across challenging scenarios like noise, imbalance, and few-shot learning.
  • We propose an expressive and generalizable GNN model, which enhances expressivity with a $k$-path rooted subgraph approach for counting paths and cycles, and improves generalization by adaptive graph contrastive learning that removes unimportant edges to handle noise and diverse domains.
  • Extensive experiments on 13 graph classification datasets, 14 graph regression datasets, and various scenarios (e.g., few-shot and noisy graphs) demonstrate the superior performance of the proposed method over state-of-the-art GNNs.
2. Preliminary and Related Work
@@
In this section, we first introduce the graph-level tasks and then summarize existing GNNs used for these tasks. Important notations are summarized in Table 1.
@@ 2.1. Graph-level Tasks
Graphs are a fundamental data structure to model relations between nodes in real-world datasets, such as social networks (Cohen, 2016; Linghu et al., 2020), chemistry graphs (Chen et al., 2024; Cohen, 2016), etc. Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\{0,1\}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively. For each node $v\in V_{i}$, $N_{i}(v)=\{u:\mathbf{A}_{i}[u][{v}]=1\}$ is the $1$-hop neighbors of ${v}$ and $N_{i}^{l}(v)$ is the neighbors of $v$ within $l$ hops. In general, there are two types of graph-level tasks, graph classification and graph regression.
These two type of tasks involve predicting either categorical labels (classification) or a continuous value (regression) for an entire graph. Formally, given a labeled graph dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, each graph $G_{i}$ is labeled with discrete label vector $y_{i}\in\{0,1\}^{\mathcal{|Y|}}$ on labels $\mathcal{Y}$ for classification, or a continuous value $y_{i}\in\mathbb{R}$ for regression. The goal is to learn a mapping function $f_{\theta}$ that generalizes to unseen graphs as follows:
(1) $f_{\theta}:G_{j}\to y_{j},\ \text{where}\ y_{j}\in\{0,1\}^{\mathcal{|Y|}}\ \text{or}\quad y_{j}\in\mathbb{R}.$
Graph classification and graph regression are critical and popular across various domains, such as databases, chemistry, and biology. Specifically, in database and data management fields, graph classification tasks include query plan strategy classification (Zhao et al., 2022a; Zhou et al., 2020) and community classification (Fang et al., 2020; Su et al., 2022), which help optimize database queries and understand data relationships. Graph regression tasks involve subgraph counting (Zhao et al., 2021; Li and Yu, 2024) and cardinality estimation of graph queries (Schwabe and Aco, 2024; Wang et al., 2022a; Teng et al., 2024). Also, biology benefits from graph classification for protein structure analysis and motif identification (Zang et al., 2023), and graph regression helps predict gene expression levels (Otal et al., 2024).
2.2. Graph Neural Networks
Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024b; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures. Existing GNNs for graph-level tasks can generally be categorized into five types: node-based, pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
2.2.1. Node-based Graph Neural Networks
As shown in Figure 1 (a), node-based GNNs first learn representations for individual nodes, which then are used to form a graph-level representation. Generally, each GNN $f_{\theta}$ consists of two basic operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be formulated as two trainable parameter matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ and a node $v\in V_{i}$, in the $l$-th layer, $\mathsf{AGG}^{(l)}(\cdot)$ in Equation (2) first aggregates the hidden representation $\mathbf{h}_{i}^{(l-1)}(u)$ and edge representation $\mathbf{e}_{uv}$ of each neighbor $u$ of node $v$ and obtain the aggregated result $\mathbf{m}^{(l)}_{i}(v)$. Then, as shown in Equation (3), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer representation $\mathbf{h}_{i}^{(l)}(v)$.
(4) $\mathbf{h}_{i}=\mathsf{READOUT}(\mathbf{H}_{i}(V_{i})),$
Model Optimization. Then, we can use a decoder (e.g., 2-layer MLPs (Zhu et al., 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$. In summary, given the labeled training dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$ and the prediction $y^{*}_{i}$ for each graph $G_{i}$ by the GNN $f_{\theta}$, the GNN $f_{\theta}$ can be optimized by minimizing the task loss as follows:
2.2.2. Hierarchical Pooling-based Graph Neural Networks
As mentioned in Equations (2) and (3), at the $l$-th layer, node-based GNNs learn node representations using the adjacency matrix $\mathbf{A}_{i}$ and the node hidden representations $\mathbf{H}_{i}^{(l-1)}(V_{i})$. In contrast, at the $l$-th layer, HP-based GNNs first generate a new cluster assignment matrix $\mathbf{S}_{i}^{(l)}\in\{0,1\}^{n_{l-1}\times n_{l}}$, where $n_{l}<n_{l-1}$. This matrix maps the nodes $V_{i}^{(l-1)}$ from the $(l-1)$-th layer to a cluster index in $\{1,\ldots,n_{l}\}$. The new adjacency matrix $\mathbf{A}_{i}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ is then reconstructed based on the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ as follows:
@@ 2.2.2. Hierarchical Pooling-based Graph Neural Networks
$y_{i}$ Discrete label or continuous property value of $G_{i}$
$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of node $v$ in $G_{i}$
$D_{i}(v),\bar{D}_{i}$ Degree of node $v$ and average degree in $G_{i}$
@@
$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer node representation of $v$ in $G_{i}$
@@
$\mathbf{H}^{(l)}_{i}(V_{j})$ The $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
@@
$imp_{v,u}$ Edge importance score between $v$ and $u$
$\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ $k$-tuple set of each node $v_{i}\in V$
$G^{s}_{\vec{t}_{v_{i}}^{k}}$ $k$-path rooted subgraph based on $\vec{t}_{v_{i}}^{k}$
$G^{s,a}_{\vec{t}_{v_{i}}^{k}}$ Feature augmented subgraph for $G^{s}_{\vec{t}_{v_{i}}^{k}}$
@@
$\mathcal{L}_{cy}(\cdot)$ Consistency-aware loss (Equation (23))

### v2

# Revisiting Graph Neural Networks for Graph-level Tasks: Taxonomy, Empirical Study, and Future Directions

@@ Abstract.
Graphs are fundamental data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems. Graph-level tasks, which involve predicting properties or labels for entire graphs, are crucial for applications like molecular property prediction and subgraph counting. While Graph Neural Networks (GNNs) have shown significant promise for these tasks, their evaluations are often limited by narrow datasets, task coverage, and inconsistent experimental setups, hindering their generalizability. In this paper, we present a comprehensive experimental study of GNNs on graph-level tasks, systematically categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs. To address these challenges, we propose a unified evaluation framework OpenGLT for graph-level GNNs. OpenGLT standardizes the evaluation process across diverse datasets, multiple graph tasks (e.g., classification and regression), and real-world scenarios, including noisy, imbalanced, and few-shot graphs. Extensive experiments are conducted on 16 baseline models across five categories, evaluated on 13 graph classification and 13 graph regression datasets. These experiments provide comprehensive insights into the strengths and weaknesses of existing GNN architectures.
Graph Neural Networks, Graph-level tasks, Unified evaluation
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 9–13, 2026; Jeju, Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’26), August 9–13, 2026, Jeju, Koreaisbn: 978-1-4503-XXXX-X/26/08copyright: none
@@ 1. Introduction
Graphs represent objects as nodes and their relationships as edges, which are key data structures across various domains to model complex interactions (Bonifati et al., 2024; Fan, 2022; Fang et al., 2022), such as social networks, molecular structures, biological systems, etc. Early graph representation learning methods (Perozzi and others, 2014; Grover and Leskovec, 2016; Wang et al., 2016; Dong et al., 2017) laid the foundation of graph learning. Graph-level tasks aim at predicting properties of an entire graph, rather than individual nodes or edges. These tasks are crucial in domains including subgraph counting in database management (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020a), and protein classification in bioinformatics (Morris et al., 2020; Li et al., 2024a). Recently, graph neural networks (GNNs) (Wang et al., 2023b; Xiang et al., 2025; Liu et al., 2025; Song et al., 2023; Liao et al., 2022; Cui et al., 2021; Li et al., 2024e) have emerged as powerful tools for graph-structured and anomaly detection tasks (Ma et al., 2023). They learn node representations by iteratively aggregating neighbor information to obtain graph representations.
Depending on the approach, we categorize GNNs for graph-level tasks into five categories: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based methods. Node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) compute node representations through message passing and aggregate them via a permutation-invariant readout function, such as averaging, to form the final graph embedding. HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) apply pooling operations to reduce the graph size and capture hierarchical structure, yielding multi-level graph representations. Subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) divide the graph into subgraphs, learn a representation for each, and then aggregate these to represent the whole graph. GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d, b) enhance graph quality by reconstructing structure and features. SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) pretrain on unlabeled data, either by predicting graph properties or maximizing agreement between augmented views of the same graph.
@@
Although various GNNs have been designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2020; Li et al., 2023; Wang et al., 2024b). To ensure fair comparison, we identify and address three key shortcomings in current evaluation frameworks.
  • Issue 1. No clear taxonomy for GNNs on graph-level Tasks. Graph-level tasks require different approaches than node-level tasks, yet a clear taxonomy for GNNs on graph-level tasks is lacking. This gap hinders holistic understanding and systematic comparison of models.
  • Issue 2. Restricted Coverage of GNN Architectures. Most evaluations focus on a limited set of architectures, such as node-based GNNs, while often ignoring more expressive models like subgraph-based GNNs. This narrow coverage limits performance comparisons and overlooks the strengths of diverse approaches.
@@
  • Issue 3. Insufficient Data Diversity. Current evaluations typically use datasets from a narrow range of domains, such as chemistry or biology. This limited diversity can lead to overfitting and restricts the generalizability of GNNs to other domains like social networks or different types of graphs.
  • Issue 4. Narrow Task and Scenario Scope. Current evaluation frameworks typically focus on a single type of graph-level task, such as molecular graph classification, and overlook diverse applications like cycle or path counting. They also assume access to ample clean labeled data, neglecting real-world challenges such as noise, class imbalance, or limited labeled graphs.
  • Issue 5. Inconsistent Evaluation Pipelines. The lack of a standardized evaluation pipeline results in inconsistent comparisons. Different works often use varying data splits, tuning protocols, and evaluation metrics, hindering fair assessment of model performance. A unified evaluation framework is needed for transparent and reliable comparisons.
@@
To address these five issues, we introduce OpenGLT, a comprehensive framework designed to provide a fair and thorough assessment of GNNs for graph-level tasks. We begin by systematically categorizing existing GNNs for graph-level tasks into five distinct types and conduct an in-depth analysis of each type to understand their unique strengths and limitations. Next, we introduce a unified evaluation pipeline with standardized data splitting and experimental settings. This pipeline incorporates graph datasets from diverse domains, including biology, chemistry, social networks, and motif graphs, ensuring broad and representative evaluations. Furthermore, we comprehensively evaluate the effectiveness and efficiency of existing graph-level GNNs on both graph classification and graph regression tasks, such as cycle counting. Beyond standard tasks, we also evaluate GNNs in real-world scenarios, including noisy graphs, imbalanced datasets, and few-shot learning settings. This unified framework ensures a holistic and equitable evaluation of GNNs, promoting fair comparisons and advancing their applicability across various graph-level tasks.
@@
We summarize the contributions of this paper on graph-level benchmarking as follows.
  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, with in-depth analysis in Section 3.
  • We propose a unified open-source evaluation framework, OpenGLT, which covers diverse tasks, datasets, and scenarios.
  • We conduct extensive experiments on 13 graph classification and 13 graph regression datasets, systematically comparing the effectiveness and efficiency of state-of-the-art GNNs, providing insights into their relative strengths, limitations, and directions.
2. Preliminaries and Related Work
@@
In this section, we first introduce graph-level tasks and then review existing experimental studies related to these graph-level tasks. Important notations are summarized in Table 1.
@@ 2.1. Graph-level Tasks
Graphs are a fundamental data structure to model relations between nodes in real-world datasets. Consequently, graph-level tasks, including classification and regression, are widely applied across domains, such as recommendation systems in social networks (Cohen, 2016; Linghu et al., 2020), molecular property and activity prediction in chemistry (Chen et al., 2024; Cohen, 2016; Mandal et al., 2022; Niazi and Mariam, 2023; Gilmer et al., 2017), and protein analysis, motif identification, or gene expression prediction in biology (Zang et al., 2023; Otal et al., 2024). Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\mathbb{R}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively. In general, graph-level tasks are to learn a mapping from $G_{i}$ to a label (classification) or a continuous value (regression).
2.2. Related Works of Experimental Studies
As summarized in Table 2, existing benchmarks and surveys (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b) often face four key limitations (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b): (i) a lack of systematic taxonomy for graph-level tasks; (ii) a predominant focus on node-based GNNs (Dwivedi et al., 2023; Hu et al., 2020a; Morris et al., 2020), leaving diverse architectures partially absent with poor categorization; (iii) the neglect of realistic scenarios such as noise, imbalance, and few-shot settings, which hinders robustness assessment; and (iv) a lack of comprehensive efficiency metrics for usability evaluation.
@@ 2.2. Related Works of Experimental Studies
$y_{i}$ Discrete label or continuous value of $G_{i}$
$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of $v$ in $G_{i}$
@@
$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer representation of $v$ in $G_{i}$
$m^{(l)}_{i}(v)$ The $l$-th layer message aggregated intermediate for node $v$
$e_{uv}$ Edge feature of edge $(u,v)$ in $G_{i}$
@@
$\mathbf{H}^{(l)}_{i}(V_{j})$ $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
@@
$\mathrm{READOUT}(\cdot)$ Graph-level aggregation function producing $\mathbf{h}_{i}$ from node/cluster embeddings
@@
$\mathcal{L}_{ssl}(\cdot)$ Self-supervised learning loss (Equation (11))
$\mathcal{L}_{cl}(\cdot)$ Contrastive learning loss (Equation (15))
Table 2. Summary of existing surveys and experimental studies on GNNs for graph-level tasks. Sur. and Exp. denote Survey and Experiments, respectively. Taxo. denotes Taxonomy, Subg. denotes Subgraph. Additionally, FewS., Imba., Effect., and Effic. denote Few-shot, Imbalanced, Effectiveness, and Efficiency, respectively.
Paper Paper Type Taxo. GNN Type Data Type Scenarios Eval Metric
Sur. Exp. Node Pool Subg. GL SSL SN BIO CHE MC Clean Noise FewS. Imba. Effect. Effic.
GNNS (Wu et al., 2020)
TUD (Morris et al., 2020)
OGB (Hu et al., 2020a)
GNNB (Dwivedi et al., 2023)
ReGCB (Li et al., 2024c)
FGNNB (Errica et al., 2020)
GPB (Wang et al., 2024b)
OpenGLT
3. Graph Neural Networks for Graph-level Tasks
Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024a; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures. Existing GNNs for graph-level tasks can generally be categorized into five types: node-based, pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
3.1. Node-based GNNs
As shown in Figure 1 (a), node-based GNNs first learn latent representations for individual nodes, which then are used to form a graph-level representation. Generally, each GNN $f_{\theta}$ consists of two fundamental operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be parameterized by learnable matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$, respectively. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ and a node $v\in V_{i}$, in the $l$-th layer, $\mathsf{AGG}^{(l)}(\cdot)$ in Equation (1) first aggregates the hidden representation $\mathbf{h}_{i}^{(l-1)}(u)$ and edge representation $\mathbf{e}_{uv}$ of each neighbor $u$ of node $v$ and obtain the aggregated result $\mathbf{m}^{(l)}_{i}(v)$. Subsequently, as shown in Equation (2), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer graph-level representation $\mathbf{h}_{i}^{(l)}(v)$.
(3) $\mathbf{h}_{i}=\mathsf{READOUT}(\mathbf{H}_{i}(V_{i})),$
Model Optimization. Then, we can use a decoder (e.g., 2-layer Multi-Layer Perceptrons (MLPs) (Zhu and others, 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$. In summary, given the labeled training dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$ and the prediction $y^{*}_{i}$ for each graph $G_{i}$ by the GNN $f_{\theta}$, the GNN $f_{\theta}$ can be optimized by minimizing the task loss as follows:
3.2. Hierarchical Pooling-based GNNs